Census ~ In 2020, Pew Research conducted a poll to examine how likely a person was to participate in the 2020 US Census based on their age. In a random sample of US adults aged 18 - 49 years old, 69.90% said they were likely to participate in the Census. The estimated standard error of the sample proportion is 0.0105.

A normal distribution can be used to model the sampling distribution of sample proportion. The researchers wish to construct a 95% confidence interval for the actual proportion of US adults aged 18 – 49 years old who said they were likely to participate in the Census.

What is the margin of error for the 95% confidence interval? Give your answer to 4 decimal places.
Answer #1

For 95% confidence interval, critical value = z = 1.96.
Margin of error = 1.96*(Standard error) = 1.96*0.0105= 0.0206